# Transonic accretion flow in the mini discs of a binary black hole system

### v1

@@ IV.3 Physical properties of the CPD with shock
In this section, we focus on studying the effect of binary parameters ($q,z_{2}$) on the shock-induced accretion solutions and examine their corresponding physical properties. In Fig. 5a, we present the dynamical structure of the shock solutions for different values of mass ratio ($q$) with binary separation $z_{2}=350$, where the Mach number ($M$) is plotted as a function of radial distance ($x$). Here, we choose $\lambda=3.45$ and $E=1.01$. The solid curves with black, blue, red, and green colors denote the shock solutions for $q=0.25,0.5,0.75$, and $1$, respectively. However, the dashed curves with the same color code represent the respective shock-free solutions. The locations of the critical points ($x_{\text{in}},x_{\text{out}}$) and shock radius ($x_{\text{sh}}$) for these shock solutions are tabulated in Table 2. In a general scenario, the accretion solution, after passing through the outer critical point $x_{\text{out}}$, continues to move toward the central black hole until it crosses the event horizon (see the dashed curves). Meanwhile, the flow slows down when the centrifugal repulsion due to flow rotation becomes dominant against the gravitational pull of the central black hole. Consequently, this supersonic flow jumps into the subsonic branch, which contains $x_{\text{in}}$. Thereafter, the flow becomes supersonic once more after crossing $x_{\text{in}}$, and continues to accrete toward the horizon. Therefore, in the shock scenario, the accretion solution can pass through both $x_{\text{in}}$ and $x_{\text{out}}$ via a discontinuous jump in between them. To calculate the shock location, we use the following relativistic shock conditions which are given by [70],
@@
    Mass flux conservation: $[\rho u^{x}]$,
@@
    Energy flux conservation: $[(e+p)u^{x}u^{t}]$,
@@
where the quantities with the square bracket denote the difference of their values across the shock front. Using these shock conditions, we calculate the shock locations at $x_{\text{sh}}=19.8524,20.9639,22.2453$, and $23.7268$ for $q=0.25,0.5,0.75$, and $1$, respectively, as indicated by the solid vertical lines. It is observed that the shock location moves away from the horizon as the mass ratio increases. In the respective panels (b)–(f) of Fig. 5, we present the profiles of radial-velocity ($v$), mass density ($\rho$), electron temperature ($T_{e}$), frequency-integrated emissivity ($\mathcal{E}$), and effective optical depth ($\tau_{\text{eff}}$) for both shock-free and shock-induced accretion solutions in Fig. 5a. In this work, we consider a SMBHB with the primary black hole mass $m_{1}=10^{6}M_{\odot}$, where $M_{\odot}$ is the Solar mass. The mass accretion rate of the flow is taken as $\dot{M}=10^{-5}\dot{M}_{\text{Edd}}$, where $\dot{M}_{\text{Edd}}=1.39\times 10^{18}m_{1}/M_{\odot}~{}\text{gm}~{}\text{s}^{-1}$ is the Eddington mass accretion rate. We observed that all the analyzed quantities significantly change at the shock locations. Moreover, the differences in these quantities across the shock fronts diminish as the mass ratio increases. This result is expected because the shock fronts settle at larger radii as the mass ratio increases. Since the radial velocity drops down at the shock radius, according to the conservation of the mass flux across the shock front (see the shock condition (a)), the mass density of the post-shock corona (hereafter PSC) increases. After the shock transition, the kinetic energy of the post-shock flow is converted into the thermal energy, leading to an increase in the electron temperature of the PSC. It is important to note that the soft photons emitted by the pre-shock disk may undergo inverse Compton scattering by the swarm of hot electrons in the PSC, resulting in the production of high-energy radiations [11]. We also observe that the emissivity of the radiation in the PSC is greater than that in the pre-shock disk due to the higher mass density and electron temperature of the flow in the PSC compared to the pre-shock disk. Furthermore, we find that the effective optical depth of the accretion disk is substantially low, even in the case of shock solutions. As a result, the disk remains optically thin, allowing the photons to escape the medium before being absorbed by the accretion disk.
@@
Now, we investigate the spectral energy distribution (SED) of the CPD by considering the emission of thermal bremsstrahlung radiation from the accretion disc. We calculate the SED for the accretion solutions presented in Figs. 5 and 6 using Eq. (22). The results are shown in Figs. 7a-b, respectively, where the variation of the quantity $\nu_{o}L_{\nu_{o}}$ with the observed frequency $\nu_{0}$ is depicted. In both panels, we observe that the emitted radiation reaches its maximum power at $\nu_{o}\approx 10^{20}\text{Hz}$ [45]. Since the electron temperature at the inner edge of the disc is $T_{e0}\approx 10^{10}\text{K}$, the spectra exhibit a cutoff at $\nu_{o}\approx 10^{21}\text{Hz}$ ($=k_{B}T_{e0}/h$) [53]. It is also found that the SEDs for the shock solutions are higher than those for the shock-free solutions [68]. This occurs because the temperature distribution in the disc for a shock solution is higher, particularly in the PSC, compared to that in a shock-free solution. Furthermore, across the entire frequency domain, the CPD spectra show negligible dependence on the binary parameters $q$ and $z_{2}$ for both the shock-induced and shock-free solutions. This is because the electron temperature of the CPD is minimally affected by $q$ and $z_{2}$ (see Figs. 5d and 6d). It is important to note that the spectral analysis of the CBD under irradiation from two mini discs was investigated in [19]. In that study, the authors demonstrated that the CBD spectrum is almost independent of the binary mass ratio. However, in our study, we obtain similar results regarding the binary mass ratio by analyzing the CPD spectrum instead of the CBD spectrum.
@@ IV.4 Shock properties of the CPD
Next, we examine the shock properties in terms of the binary parameters $q$ and $z_{2}$. To do this, we calculate the shock radius ($r_{\text{sh}}$) as a function of $q$ for different values of $z_{2}$. The obtained results are presented in Fig. 8a, where the solid (black), dashed (blue), dotted (red), and dash-dotted (green) curves correspond to $z_{2}=250$, $500$, $750$, and $1000$, respectively. In this case, the input parameters are chosen as $\lambda=3.5$ and $E=1.01$. We observe that the shock radius recedes away from the black hole as $q$ increases, irrespective of the value of $z_{2}$. Moreover, the shock disappears when $q$ exceeds a critical value, as the shock conditions are no longer satisfied beyond that limit. Also, for a given $q$, the shock radius decreases when $z_{2}$ increases. Since both the density and temperature of the flow undergo significant changes at the shock location, it is essential to study their variations across the shock locations. To examine the density compression, we define a quantity, which is called the compression ratio as $R=\Sigma_{+}/\Sigma_{-}$, where ``+"``"``+"` ` + " and ``"``"``-"` ` - " denote the post-shock and pre-shock values of the surface density $\Sigma$ ($=2\rho H$). Following Eq. (10), we find an expression of $R$ in terms of flow radial 3-velocity ($v$) and its Lorentz factor ($\gamma_{v}$) as $R=(v\gamma_{v})_{-}/(v\gamma_{v})_{+}$. In Fig. 8b, we present the variation of $R$ with $q$ for the same set of input parameters used in Fig. 8a. It is observed that $R$ decreases with $q$ for a given value of $z_{2}$, which is expected, as the shock locations shift to larger radii when $q$ increases. Since the Mach number ($M$) depends on the flow temperature ($\Theta$), the temperature jump at the shock front is characterized by a quantity called the shock strength ($S$). It is defined as the ratio of the pre-shock Mach number ($M_{-}$) to the post-shock Mach number ($M_{+}$) as $S=M_{-}/M_{+}$. In Fig. 8c, we show the variation of $S$ with $q$. As the shock radius increases with $q$, it is again expected that $S$ decreases as $q$ increases. Furthermore, for a fixed $q$, both $R$ and $S$ increase with $z_{2}$ as $r_{\text{sh}}$ shrinks with $z_{2}$. It is noted that the dependencies of the shock location, density compression, and temperature compression on both $q$ and $z_{2}$ are consistent with the results presented in Figs. 5 and 6.
@@ IV.5 $E-\lambda$ parameter space for shock solutions
Figure 9: Modification of the shock parameter space in the specific angular momentum ($\lambda$) and energy ($E$) plane for the mass ratios $q=0.01$, $0.1$, and $1$. In this figure, we choose $z_{2}=300$. See the text for details.
Figure 10: Modification of the shock parameter space in the $\lambda-E$ plane for the binary separations $z_{2}=250$, $500$, and $750$. In this figure, we choose $q=0.5$. See the text for details.
We mention earlier that the A-type solutions can harbor shock transitions, provided the relativistic shock conditions are satisfied. Most importantly, shock solutions are not unique but rather have an energy range for a given angular momentum (see Section IV.1). Here, we explore the available parameter space for shock solutions in the CPD and see the modification of those identified regions with the mass ratio ($q$) and binary separation ($z_{2}$). For this, we calculate an effective region of the parameter space in the specific angular momentum ($\lambda$) and energy ($E$) plane for a given combination of $q$ and $z_{2}$ that admits shock solutions. In Fig. 9, we present the modification of the shock parameter space for different values of $q$ with a fixed $z_{2}=300$. The regions bounded by the solid (blue), dotted (red), and dashed (gray) curves correspond to the results for $q=0.01$, $0.1$, and $1$, respectively. We observed that, the parameter space shifts toward lower $\lambda$ and lower $E$ domains as $q$ increases. Moreover, the area under the parameter space gradually decreases when the mass of the secondary black hole becomes comparable to that of the primary black hole. Similarly, in Fig. 10, we show the modification of the shock parameter space in the $\lambda-E$ plane as a function of $z_{2}$ with a fixed $q=0.5$, where the effective regions bounded by the solid (blue), dotted (red), and dashed (gray) are obtained for $z_{2}=250$, $500$, and $750$, respectively. It is observed that as $z_{2}$ increases, the parameter space shifts toward higher $\lambda$ and lower $E$ sides. In addition, the shock parameter space expands as the secondary black hole moves farther away from the primary one.
@@ V Conclusions
We find that flow may possess single or multiple critical points depending on the specific angular momentum ($\lambda$), energy ($E$), binary mass ratio ($q$), and binary separation ($z_{2}$). We obtain the A-type accretion solutions that contain multiple critical points and show that the behaviors of the accretion solution solutions remain A-type with an increase of both $q$ and $z_{2}$. Since the A-type accretion solutions can exhibit standing shock transitions, provided the relativistic shock conditions are satisfied, we also investigate the shock-induced accretion solutions. We analyze the shock properties such as shock radius ($r_{\text{sh}}$), compression ratio ($R$), and shock strength ($S$) as functions of $q$ and $z_{2}$. We observe that as $q$ increases, both $R$ and $S$ decreases as $r_{\text{sh}}$ moves away from the horizon. However, as $z_{2}$ increases, $r_{\text{sh}}$ decreases, resulting in higher values of $R$ and $S$. We further investigate the luminosity distribution of the CPD as a function of $q$ and $z_{2}$. We find that the CPD spectra barely depend on $q$ and $z_{2}$ for both the shock-induced and shock-free solutions, which agrees with the findings in [19]. Also, we examine the shock parameter space in the $\lambda-E$ plane and observe how it is modified by $q$ and $z_{2}$. We find that as $q$ increases, the shock parameter space shrinks, whereas it expands with an increase in $z_{2}$. Therefore, the possibility of shock transitions is greater for smaller mass ratios and larger binary separations. Indeed, this analysis implies that while the shock properties of the CPD moderately change due to the presence of a secondary black hole, it does not significantly affect the CPD spectrum. As mentioned earlier, these results yield similar consequences of the binary mass ratio, as demonstrated in [19]. The previous study [19] analyzed the CBD accretion and its associated spectrum in the presence of irradiation from CPD and CSD, treating all disc components as standard discs within a Newtonian framework rather than considering general relativity effects. However, in our study, we focus solely on the accretion properties of the CPD within a fully general relativistic framework. Additionally, we consider the CPD as an optically thin medium that does not emit like a black body, as is typically assumed in standard disc models. Therefore, even though our analysis differs from the work in [19], the remarks on the binary mass ratio are the same.
@@
Finally, we mention the limitations of our work. We do not incorporate viscosity [65, 68], magnetic fields [71, 67, 72], thermal conduction [73, 74], and radiative cooling [54, 58] in the flow equations. Moreover, we only consider the thermal bremsstrahlung emission, neglecting the synchrotron and Compton emission processes [75, 58]. However, their presence in an accretion model is expected to provide more valuable insights into accretion physics. Also, we consider a binary system consisting of two Schwarzschild black holes. In reality, astrophysical black holes possess some spin. Exact analytical solutions for a stable binary configuration of two spinning black holes are already available in the literature [76]. We expect, the spin parameters may also effect the CPD accretion flow. We plan to investigate all these aspects in future work and report our findings elsewhere.

### v2

@@ IV.3 Physical properties of the CPD with shock
In this section, we focus on studying the effect of binary parameters ($q,z_{2}$) on the shock-induced accretion solutions and examine their corresponding physical properties. In Fig. 5a, we present the dynamical structure of the shock solutions for different values of mass ratio ($q$) with binary separation $z_{2}=350$, where the Mach number ($M$) is plotted as a function of radial distance ($x$). Here, we choose $\lambda=3.45$ and $E=1.01$. The solid curves with black, blue, red, and green colors denote the shock solutions for $q=0.25,0.5,0.75$, and $1$, respectively. However, the dashed curves with the same color code represent the respective shock-free solutions. The locations of the critical points ($x_{\text{in}},x_{\text{out}}$) and shock radius ($x_{\text{sh}}$) for these shock solutions are tabulated in Table 2. In a general scenario, the accretion solution, after passing through the outer critical point $x_{\text{out}}$, continues to move toward the central black hole until it crosses the event horizon (see the dashed curves). Meanwhile, the flow slows down when the centrifugal repulsion due to flow rotation becomes dominant against the gravitational pull of the central black hole. Consequently, this supersonic flow jumps into the subsonic branch, which contains $x_{\text{in}}$. Thereafter, the flow becomes supersonic once more after crossing $x_{\text{in}}$, and continues to accrete toward the horizon. Therefore, in the shock scenario, the accretion solution can pass through both $x_{\text{in}}$ and $x_{\text{out}}$ via a discontinuous jump in between them. To calculate the shock location, we use the following relativistic shock conditions which are given by Taub (1948),
@@
    Mass flux conservation: $\rho u^{x}]$,
@@
    Energy flux conservation: $(e+p)u^{x}u^{t}]$,
@@
where the quantities with the square bracket denote the difference of their values across the shock front. Using these shock conditions, we calculate the shock locations at $x_{\text{sh}}=19.8524,20.9639,22.2453$, and $23.7268$ for $q=0.25,0.5,0.75$, and $1$, respectively, as indicated by the solid vertical lines. It is observed that the shock location moves away from the horizon as the mass ratio increases. In the respective panels (b)–(f) of Fig. 5, we present the profiles of radial-velocity ($v$), mass density ($\rho$), electron temperature ($T_{e}$), frequency-integrated emissivity ($\mathcal{E}$), and effective optical depth ($\tau_{\text{eff}}$) for both shock-free and shock-induced accretion solutions in Fig. 5a. In this work, we consider a SMBHB with the primary black hole mass $m_{1}=10^{6}M_{\odot}$, where $M_{\odot}$ is the Solar mass. The mass accretion rate of the flow is taken as $\dot{M}=10^{-5}\dot{M}_{\text{Edd}}$, where $\dot{M}_{\text{Edd}}=1.39\times 10^{18}m_{1}/M_{\odot}~\text{gm}~\text{s}^{-1}$ is the Eddington mass accretion rate. We observed that all the analyzed quantities significantly change at the shock locations. Moreover, the differences in these quantities across the shock fronts diminish as the mass ratio increases. This result is expected because the shock fronts settle at larger radii as the mass ratio increases. Since the radial velocity drops down at the shock radius, according to the conservation of the mass flux across the shock front (see the shock condition (a)), the mass density of the post-shock corona (hereafter PSC) increases. Note that the mass density lies in the range of approximately $10^{-16}-10^{-14}$ gm cm-3. Since we have adopted a very small value of $\dot{M}$, as mentioned above, Eq. (12) yields correspondingly low values of $\rho$. This choice of $\dot{M}$ is reasonable, as we are dealing with a radiatively inefficient accretion flow (RIAF), where $\dot{M}<<\dot{M}_{\rm Edd}$. For example, in the case of Sgr A*, $\dot{M}\simeq 4.5\times 10^{-7}\dot{M}_{\rm Edd}$, and the corresponding $\rho$ is estimated to lie in the range $10^{-18}-10^{-14}$ gm cm-3 Yuan et al. (2003). Another example is M87*, for which $\dot{M}\lesssim 9.2\times 10^{-4}\dot{M}_{\rm Edd}$ Kuo et al. (2014). Therefore, also for M87*, one might expect $\rho$ to lie within a similar range. It is worth noting that if higher values of $\dot{M}$ are considered, as suggested by some studies in the literature for Sgr A* and M87* Mościbrodzka et al. (2009); Yarza et al. (2020), then, according to Eq. (12), the corresponding values of $\rho$ will increase accordingly. Since these supermassive black hole systems exhibit very low mass densities, they are optically thin, and their luminosities are low – an important characteristic of RIAFs or hot accretion flows (HAFs) (see the review paper Yuan and Narayan (2014)). After the shock transition, the kinetic energy of the post-shock flow is converted into the thermal energy, leading to an increase in the electron temperature of the PSC. It is important to note that the soft photons emitted by the pre-shock disk may undergo inverse Compton scattering by the swarm of hot electrons in the PSC, resulting in the production of high-energy radiations Roedig et al. (2014). We also observe that the emissivity of the radiation in the PSC is greater than that in the pre-shock disk due to the higher mass density and electron temperature of the flow in the PSC compared to the pre-shock disk. Furthermore, we find that the effective optical depth of the accretion disk is substantially low, even in the case of shock solutions. As a result, the disk remains optically thin, allowing the photons to escape the medium before being absorbed by the accretion disk.
@@
Now, we investigate the spectral energy distribution (SED) of the CPD by considering the emission of thermal bremsstrahlung radiation from the accretion disc. We calculate the SED for the accretion solutions presented in Figs. 5 and 6 using Eq. (24). The results are shown in Figs. 7a-b, respectively, where the variation of the quantity $\nu_{o}L_{\nu_{o}}$ with the observed frequency $\nu_{0}$ is depicted. In both panels, we observe that the emitted radiation reaches its maximum power at $\nu_{o}\approx 10^{20}\text{Hz}$ Yarza et al. (2020). Since the electron temperature at the inner edge of the disc is $T_{e0}\approx 10^{10}\text{K}$, the spectra exhibit a cutoff at $\nu_{o}\approx 10^{21}\text{Hz}$ ($k_{B}T_{e0}/h$) Novikov and Thorne (1973). It is also found that the SEDs for the shock solutions are higher than those for the shock-free solutions Patra et al. (2024b). This occurs because the temperature distribution in the disc for a shock solution is higher, particularly in the PSC, compared to that in a shock-free solution. Furthermore, across the entire frequency domain, the CPD spectra show negligible dependence on the binary parameters $q$ and $z_{2}$ for both the shock-induced and shock-free solutions. This is because the electron temperature of the CPD is minimally affected by $q$ and $z_{2}$ (see Figs. 5d and 6d).
@@ IV.4 Shock properties of the CPD
Next, we examine the shock properties in terms of the binary parameters $q$ and $z_{2}$. To do this, we calculate the shock radius ($r_{\text{sh}}$) as a function of $q$ for different values of $z_{2}$. The obtained results are presented in Fig. 8a, where the solid (black), dashed (blue), dotted (red), and dash-dotted (green) curves correspond to $z_{2}=250$, $500$, $750$, and $1000$, respectively. In this case, the input parameters are chosen as $\lambda=3.5$ and $E=1.01$. We observe that the shock radius recedes away from the black hole as $q$ increases, irrespective of the value of $z_{2}$. Moreover, the shock disappears when $q$ exceeds a critical value, as the shock conditions are no longer satisfied beyond that limit. Also, for a given $q$, the shock radius decreases when $z_{2}$ increases. Since both the density and temperature of the flow undergo significant changes at the shock location, it is essential to study their variations across the shock locations. To examine the density compression, we define a quantity, which is called the compression ratio as $R=\Sigma_{+}/\Sigma_{-}$, where +``+" and ``-" denote the post-shock and pre-shock values of the surface density $\Sigma$ ($2\rho H$). Following Eq. (12), we find an expression of $R$ in terms of flow radial 3-velocity ($v$) and its Lorentz factor ($\gamma_{v}$) as $R=(v\gamma_{v})_{-}/(v\gamma_{v})_{+}$. In Fig. 8b, we present the variation of $R$ with $q$ for the same set of input parameters used in Fig. 8a. It is observed that $R$ decreases with $q$ for a given value of $z_{2}$, which is expected, as the shock locations shift to larger radii when $q$ increases. Since the Mach number ($M$) depends on the flow temperature ($\Theta$), the temperature jump at the shock front is characterized by a quantity called the shock strength ($S$). It is defined as the ratio of the pre-shock Mach number ($M_{-}$) to the post-shock Mach number ($M_{+}$) as $S=M_{-}/M_{+}$. In Fig. 8c, we show the variation of $S$ with $q$. As the shock radius increases with $q$, it is again expected that $S$ decreases as $q$ increases. Furthermore, for a fixed $q$, both $R$ and $S$ increase with $z_{2}$ as $r_{\text{sh}}$ shrinks with $z_{2}$. It is noted that the dependencies of the shock location, density compression, and temperature compression on both $q$ and $z_{2}$ are consistent with the results presented in Figs. 5 and 6.
@@ IV.5 $E-\lambda$ parameter space for shock solutions
Figure 9: Modification of the shock parameter space in the specific angular momentum ($\lambda$) and energy ($E$) plane for the mass ratios $q=0.01$, $0.1$, and $1$ with binary separation $z_{2}=300$ (panel (a)), and for $z_{2}=250$, $500$, and $750$ with $q=0.5$ (panel (b)). See the text for details.
We mention earlier that the A-type solutions can harbor shock transitions, provided the relativistic shock conditions are satisfied. Most importantly, shock solutions are not unique but rather have an energy range for a given angular momentum (see Section IV.1). Here, we explore the available parameter space for shock solutions in the CPD and see the modification of those identified regions with the mass ratio ($q$) and binary separation ($z_{2}$). For this, we calculate an effective region of the parameter space in the specific angular momentum ($\lambda$) and energy ($E$) plane for a given combination of $q$ and $z_{2}$ that admits shock solutions. In Fig. 9a, we present the modification of the shock parameter space for different values of $q$ with a fixed $z_{2}=300$. The regions bounded by the solid (blue), dotted (red), and dashed (gray) curves correspond to the results for $q=0.01$, $0.1$, and $1$, respectively. We observed that, the parameter space shifts toward lower $\lambda$ and lower $E$ domains as $q$ increases. Moreover, the area under the parameter space gradually decreases when the mass of the secondary black hole becomes comparable to that of the primary black hole. Similarly, in Fig. 9b, we show the modification of the shock parameter space in the $\lambda-E$ plane as a function of $z_{2}$ with a fixed $q=0.5$, where the effective regions bounded by the solid (blue), dotted (red), and dashed (gray) are obtained for $z_{2}=250$, $500$, and $750$, respectively. It is observed that as $z_{2}$ increases, the parameter space shifts toward higher $\lambda$ and lower $E$ sides. In addition, the shock parameter space expands as the secondary black hole moves farther away from the primary one.
@@ V Conclusions
We find that flow may possess single or multiple critical points depending on the specific angular momentum ($\lambda$), energy ($E$), binary mass ratio ($q$), and binary separation ($z_{2}$). We obtain the A-type accretion solutions that contain multiple critical points and show that the behaviors of the accretion solution solutions remain A-type with an increase of both $q$ and $z_{2}$. Since the A-type accretion solutions can exhibit standing shock transitions, provided the relativistic shock conditions are satisfied, we also investigate the shock-induced accretion solutions. We analyze the shock properties such as shock radius ($r_{\text{sh}}$), compression ratio ($R$), and shock strength ($S$) as functions of $q$ and $z_{2}$. We observe that as $q$ increases, both $R$ and $S$ decreases as $r_{\text{sh}}$ moves away from the horizon. However, as $z_{2}$ increases, $r_{\text{sh}}$ decreases, resulting in higher values of $R$ and $S$. We further investigate the luminosity distribution of the CPD as a function of $q$ and $z_{2}$. We find that the CPD spectra barely depend on $q$ and $z_{2}$ for both the shock-induced and shock-free solutions, which agrees with the findings in Lee et al. (2024). Also, we examine the shock parameter space in the $\lambda-E$ plane and observe how it is modified by $q$ and $z_{2}$. We find that as $q$ increases, the shock parameter space shrinks, whereas it expands with an increase in $z_{2}$. Therefore, the possibility of shock transitions is greater for smaller mass ratios and larger binary separations. Indeed, this analysis implies that while the shock properties of the CPD moderately change due to the presence of a secondary black hole, it does not significantly affect the CPD spectrum.
It is worth noting that a recent study by Lee et al. (2024) investigated the spectral features of BBH accretion flows. The authors analyzed CBD accretion and its associated spectrum in the presence of irradiation from CPD and CSD, treating all disc components as standard discs within a Newtonian framework. Additionally, the accretion discs were assumed to be coplanar. In their study, the authors showed that the CBD spectrum is nearly independent of the binary mass ratio. However, in our study, we focus solely on the accretion properties of the CPD within a fully general relativistic framework. Since the present symmetry of the BBH metric (Eq. (1)) allows the fluid to rotate around the Z-axis to conserve angular momentum, the accretion discs of the two black holes would not be coplanar in this scenario. Additionally, we consider the CPD as an optically thin medium that does not emit like a black body, as is typically assumed in standard disc models. Therefore, although our analysis, including the disc configuration, differs entirely from that of Lee et al. (2024), the observations regarding the binary mass ratio are qualitatively similar.
@@
Finally, we mention the limitations of our work. We do not incorporate viscosity Dihingia et al. (2019); Patra et al. (2024b), magnetic fields Mitra et al. (2022); Mitra and Das (2024); Dihingia and Mizuno (2024), thermal conduction Mitra et al. (2023); Singh and Das (2024), and radiative cooling Sarkar and Chattopadhyay (2018, 2022) in the flow equations. Moreover, we only consider the thermal bremsstrahlung emission, neglecting the synchrotron and Compton emission processes Dihingia et al. (2020a); Sarkar and Chattopadhyay (2022). However, their presence in an accretion model is expected to provide more valuable insights into accretion physics. Also, we consider a binary system consisting of two Schwarzschild black holes. In reality, astrophysical black holes possess some spin. Exact analytical solutions for a stable binary configuration of two spinning black holes are already available in the literature Astorino and Viganò (2022). We expect, the spin parameters may also effect the CPD accretion flow. We plan to investigate all these aspects in future work and report our findings elsewhere.